TheSmartOne (thesmartone):

Helped by: Michele_Laino , amistre64 Find the equivalent algebraic expression for the composition below. \(\sf cos(arctan(x))=?\) (Simplify your answer.)

TheSmartOne (thesmartone):

@amistre64 @Michele_Laino

OpenStudy (anonymous):

Find an equivalent algebraic expression for each composition.?

OpenStudy (anonymous):

here this mite help @TheSmartOne For part A: Draw a right triangle. Inverse cos(x) is really telling you that the adjacent side and the hypotenuse are x and 1 respectively. Label these on the triangle. Using the pythagorean theorem, u find that the opposite side is the sqrt(1-x^2). So sin(arccos(x)) is sqrt(1-x^2). Using the same logic you should get 1/(sqrt(x^2+1)) for cos(arctan(

TheSmartOne (thesmartone):

The answer is in terms of x. And that is what the question said exactly.

OpenStudy (amistre64):

im with love on it ....

TheSmartOne (thesmartone):

The example says to make \(\sf arctan(x)=\theta\)

TheSmartOne (thesmartone):

so then \(\sf tan\theta =x\)

OpenStudy (anonymous):

sorry is sopost to say Using the same logic you should get 1/(sqrt(x^2+1)) for cos(arctan(x0

TheSmartOne (thesmartone):

And then using that information to draw the right angle triangle.

OpenStudy (anonymous):

right

OpenStudy (amistre64):



OpenStudy (anonymous):

thanks for the medal

TheSmartOne (thesmartone):

And then \(\sf\Large cos\theta = \frac{adj}{hyp}=\frac{1}{\sqrt{x^2+1}}\)

TheSmartOne (thesmartone):

Now it all makes sense. Thanks a lot! @Loveheart @amistre64 :)

OpenStudy (anonymous):

@TheSmartOne do you need help with any thing els

TheSmartOne (thesmartone):

Not at the moment. :)

OpenStudy (anonymous):

kk just tag me when you do

OpenStudy (michele_laino):

let's suppose that \[\theta = \arctan x\] then we have: \[\tan \theta = x\]

OpenStudy (michele_laino):

so we can write: \[\sin \theta = x\cos \theta \]

OpenStudy (michele_laino):

now I use the fundamental identity, and I get: \[{\left( {\sin \theta } \right)^2} + {\left( {\cos \theta } \right)^2} = 1\]

OpenStudy (michele_laino):

after a substitution, I can write: \[{\left( {x\cos \theta } \right)^2} + {\left( {\cos \theta } \right)^2} = 1\]

OpenStudy (michele_laino):

and finally: \[\cos \theta = \pm \frac{1}{{\sqrt {1 + {x^2}} }}\]

TheSmartOne (thesmartone):

Interesting way to arrive at the same answer. :)

OpenStudy (amistre64):

i stretching ropes is quicker lol

OpenStudy (amistre64):

thats just the ancient egyptian in me tho .... all these new fangled maths are just hogwash ...

TheSmartOne (thesmartone):

Well, I like Michele's method. Saves time because you don't have to graph. While the book said to graph it so that maybe you can have a visual aspect to it. Thank you all! :D

OpenStudy (michele_laino):

thank you! @TheSmartOne @amistre64

eclipsedstar (eclipsedstar):

@loveheart don't forget to cite your sources, add that you got that from yahoo.

eclipsedstar (eclipsedstar):

http://prntscr.com/6qj7av